The equation for voltage dividers in electronics stated in basic algebraic terms is:
x=y(b/(a+b))
As a and b represent resistor values I have the equations to solve for each of them. I have been able to figure out the steps to calculate for a, however, the steps to arrive at the formula for b has eluded me. The formula given is:
b=(xa)/(y-x)
I suppose I could just memorize the formula but I think I must be missing something very basic that I cannot work out the steps myself. Your help would be much appreciated. Glenn

Answer by rothauserc(4718)

we need to solve
x=y(b/(a+b)) for b
divide both sides of = by y
x/y = b/(a+b)
cross multiply the two fractions and we get
x(a+b) = yb
xa+xb = yb
now group the terms with b on to same side of = sign
yb-xb = xa
factor out the b
b(y-x) = xa and
b = xa/(y-x)